A sample of gas expands and increases its volume by 2.5 L while pushing a piston against an external force of 2 kPa.

 

What is the work done by the expanding gas?


1
Expert's answer
2021-10-05T01:46:02-0400

"Pressure , P=2KPa=2000Pa=2000atm"

change in volume, "\Delta V=2.5L"

work done, "W=-P\Delta V=-2000atm\times 2.5L=-5000L.atm=\frac{-5000L. atm}{1}\times \frac{101.3J}{1L.atm}=-506500J"